# “There Has To Be a Lot That We’re Missing”: Moderating AI-Generated Content on Reddit

### v6

@@ 5.2.2. Equitably Enforcing AIGC Bans
Our findings detail some of the challenges that moderators currently experience enforcing AIGC restrictions and bans. Even before the AIGC challenge, moderators were stretched thin trying to protect their communities from problematic content (e.g. mis- and disinformation) and behavior (e.g. harassment, incivility, and influence campaigns). With detection of AIGC being an unsolved problem (edwards23), moderators rely on heuristics to enforce restrictions or bans. Our findings show that these heuristics can be time-intensive. Even worse, they may be biased, as other work has shown that people’s intuitions about AIGC are wrong in predictable ways (Jakesch_2023) and that certain social groups are more often suspected of AI use (kadoma2024generativeaiperceptualharms). Such consistent bias could disproportionately affect minorities and individuals seeking legitimate participation (liang23), such as non-native speakers, newcomers, and gender or racial minorities. It is concerning that most of our participants were confident in the efficacy of their heuristics and unaware of any potential bias. Our participants’ reliance on these heuristics suggests that moderators should be educated about potential bias and further studies are needed to evaluate the efficacy of detection heuristics in content moderation. Platforms may be able to help moderators by adding tools for AIGC detection. Earlier work demonstrated that “synthesized social signals,” or “social signals computationally derived from an account’s history,” help community members evaluate the credibility of other accounts (Im20). Similar tools may be useful for detecting AIGC. For example, showing information about a poster’s interactions with the system, such as time spent editing a post, could be helpful, though this also has user privacy implications. Surfacing a poster’s past behavior and other community memberships could also be relevant, as these have been shown to be useful in identifying misinformation posters (Bozarth2023; diakopoulos_diamonds_2010).
@@
Given the combination of concerns about AIGC and the lack of tools to detect it, we expect to see communities develop their own, non-technical, means for demonstrating authenticity. For example, past research into anonymous online communities (bernstein14) found several such practices, such as triforcing (posting complex Unicode symbols only meaningful to the community) and timestamping (posting photos of oneself holding a piece of paper with the current date and time on it). These techniques signal in-group membership and authenticity by presenting something that only an authentic poster can produce. These forms of effortful communication (Zhang22) signal the poster’s effort, which makes them seem more valuable, given the relationship between effort and value mentioned by a participant in Section 4.1.1. In the age of low-effort AIGC we expect to see online communities place an even greater premium on effortful communication, which platform design can encourage. While these social innovations may emerge naturally from within a community, platforms can provide technical affordances to support their adoption, and the HCI research community can help.
@@ 6. Limitations and Future Work
Our qualitative study is based on interviews with moderators of communities that are concerned with AIGC use and are taking steps to curb it. This recruiting criteria supports our goal of understanding some of the attitudes and practical considerations of Reddit moderators who are adapting to AIGC. We note, however, that our study and results are not meant to provide a generalizable view of how all Reddit moderators and communities consider AIGC.
Our study purposefully omitted those communities that have fully embraced AIGC, as well as those that have not yet created rules governing its use. Our recruitment additionally focused on popular subreddits in order to target those that are currently most affected by AIGC. However, our participants often also moderated smaller subreddits and were able to share how their experiences and concerns are relevant to communities of various sizes. Additional studies could include moderators from a broader set of communities to provide a more complete picture of the range of ways that online communities are responding to AIGC, and to further explore the relationship between a community’s values (Prinster24) and its stance on AIGC. We hope that future quantitative work will complement our own, by investigating a wider range of communities and making stronger inferences about the platform as a whole. For example, future work could use surveys or computational content analysis to test how our qualitative findings generalize across subreddits of different size and type. Further, since AIGC is an internet-scale phenomenon, future studies should explore how moderators are responding on other platforms, which may have different purposes and norms.
@@ 7. Conclusion
We performed a qualitative investigation into volunteer Reddit moderators’ experiences with and attitudes towards AI-generated content. To better understand how CSCW researchers can support community self-determination we focused on communities that have restricted the use of AIGC. We heard from moderators that they are concerned about AIGC impacting the content quality, social dynamics, and governance processes in their communities, despite also seeing its potential to make it easier for some community members to contribute. In response, communities have enacted bans, which moderators enforce using time-intensive and imperfect heuristics. Despite enforcement challenges, our moderators report that, for now, their rules suffice, as they serve to clarify community norms about AIGC use.

### v7

@@ 5.2.2. Equitably Enforcing AIGC Bans
Our findings detail some of the challenges that moderators currently experience enforcing AIGC restrictions and bans. Even before the AIGC challenge, moderators were stretched thin trying to protect their communities from problematic content (e.g., mis- and disinformation) and behavior (e.g., harassment, incivility, and influence campaigns). With detection of AIGC being an unsolved problem (Edwards, 2023), moderators rely on heuristics to enforce restrictions or bans. Our findings show that these heuristics can be time-intensive. Even worse, they may be biased, as other work has shown that people’s intuitions about AIGC are wrong in predictable ways (Jakesch et al., 2023) and that certain social groups are more often suspected of AI use (Kadoma et al., 2025). Such consistent bias could disproportionately affect minorities and individuals seeking legitimate participation (Liang et al., 2023), such as non-native speakers, newcomers, and gender or racial minorities. It is concerning that most of our participants were confident in the efficacy of their heuristics and unaware of any potential bias. Our participants’ reliance on these heuristics suggests that moderators should be educated about potential bias and that further studies should evaluate the efficacy of detection heuristics in content moderation. Platforms may be able to help moderators by adding tools for AIGC detection. Earlier work demonstrated that “synthesized social signals,” or “social signals computationally derived from an account’s history,” help community members evaluate the credibility of other accounts (Im et al., 2020). Similar tools may be useful for detecting AIGC. For example, showing information about a poster’s interactions with the system, such as time spent editing a post, could be helpful, though this also has user privacy implications. Surfacing a poster’s past behavior and other community memberships could also be relevant, as these have been shown to be useful in identifying misinformation posters (Bozarth et al., 2023).
@@
Given the combination of concerns about AIGC and the lack of tools to detect it, we expect to see communities develop their own, non-technical, means for demonstrating authenticity. Communities may have no choice: they may struggle to function, as recent work has suggested that AI makes cooperation harder in pseudonymous, low-trust settings like online communities (Wojtowicz and DeDeo, 2025). Past research provides examples where novel communication practices have emerged in anonymous online communities, such as triforcing (posting complex Unicode symbols only meaningful to the community) and timestamping (posting photos of oneself holding a piece of paper with the current date and time on it) (Bernstein et al., 2021). These techniques signal in-group membership and authenticity by presenting something that only an authentic poster can produce. These techniques also demonstrate effort, which—as a participant mentioned in Section 4.1.1—makes a contribution seem more valuable. Past research calls this type of communication effortful communication and documents how it can facilitate connection between online community members (Zhang et al., 2022). Certain interface designs can encourage such communication, but it remains to be seen if these techniques can offset the negative effects that AIGC may have on interpersonal communication. In the age of low-effort AIGC we expect to see online communities place an even greater premium on effortful communication, which platform design can encourage. While these social innovations may emerge naturally from within a community, platforms can provide technical affordances to support their adoption, and the HCI research community can help.
@@ 6. Limitations and Future Work
Our qualitative study is based on interviews with moderators of communities that were taking steps to govern the use of AIGC. This recruiting criteria supported our goal of understanding some of the attitudes and practices of Reddit moderators as they adapt to the growing prevalence of AIGC. Our recruitment criteria and limited sample means that our study and results do not necessarily explain how all Reddit moderators and communities consider AIGC. Fully understanding the perceptions, attitudes, and impact of generative AI on online communities requires future work that approaches this topic from different angles.
For example, our study purposefully omitted communities that have fully embraced AIGC, as well as those without AI-specific rules. Studying communities who encourage the use of AIGC could be a fruitful direction for future work. Our recruitment additionally focused on larger subreddits in order to target those that are currently most affected by AIGC. While our participants also moderated smaller subreddits and shared how their experiences and concerns are relevant to communities of various sizes, additional studies could include moderators from a wider set of communities. Such studies could provide a more complete picture of the range of ways that online communities are responding to AIGC, and to further explore the relationship between a community’s values (Prinster et al., 2024) and its stance on AIGC. For example, the AI concerns that we identified may be specific to certain subreddit topics, and our limited sample may have missed other important concerns. A recent study of ours used computational content analysis to understand how these qualitative findings generalize across subreddits of different size and type (Lloyd et al., 2025). Other work could use surveys or other quantitative methods to more fully understand the impact on communities. Further, since AIGC is an internet-scale phenomenon, future studies should look beyond Reddit in order to understand how moderators are responding on platforms with different purposes and norms.
@@ 7. Conclusion
We performed a qualitative investigation into volunteer Reddit moderators’ experiences with, and attitudes towards, AI-generated content. Moderators shared that they are concerned about AIGC impacting the content quality, social dynamics, and governance processes in their communities, despite also seeing its potential to make it easier for some community members to contribute. In response, communities have enacted bans, which moderators enforce using time-intensive and imperfect heuristics. Despite enforcement challenges, our participants reported that their rules are currently sufficient, as they clarify community norms about AIGC use.